# Stereo Hand-Object Reconstruction for Human-to-Robot Handover

### v2

@@ Abstract
Jointly estimating hand and object shape facilitates the grasping task in human-to-robot handovers. However, relying on hand-crafted prior knowledge about the geometric structure of the object fails when generalising to unseen objects, and depth sensors fail to detect transparent objects such as drinking glasses. In this work, we propose a stereo-based method for hand-object reconstruction that combines single-view reconstructions probabilistically to form a coherent stereo reconstruction. We learn 3D shape priors from a large synthetic hand-object dataset to ensure that our method is generalisable, and use RGB inputs to better capture transparent objects. We show that our method reduces the object Chamfer distance compared to existing RGB based hand-object reconstruction methods on single view and stereo settings. We process the reconstructed hand-object shape with a projection-based outlier removal step and use the output to guide a human-to-robot handover pipeline with wide-baseline stereo RGB cameras. Our hand-object reconstruction enables a robot to successfully receive a diverse range of household objects from the human.
@@ I Introduction
As humans and robots share the same workspace [1], exchanging objects is one of the essential skills to enable smooth collaboration. We consider the scenario where a human hands over various everyday objects to a robot and the robot aims to grasp the object while avoiding the human (human-to-robot handover [2, 3, 4]), and place the object at a predefined delivery location (Fig. 1). Objects can vary in their physical properties, such as appearance, transparency, shape, size, and mass [5]. Container-like objects, such as cups, drinking glasses, or boxes, can be empty or filled with content, affecting the perception of the appearance or the weight of the object [4, 6]. In this highly interactive scenario, hand-object shape reconstruction is an essential task as the object shape informs the robot about how to grasp the object, while the hand shape allows the robot to avoid dangerous contact with the human.
@@
Visual modalities (RGB and depth images) are used in hand-object shape reconstruction for human-to-robot handover. Previous works [7, 3, 8, 9, 10] used depth sensors to capture the pointcloud of the scene and applied segmentation models to extract the pointcloud of the hand and object. However, transparent objects are not easy to be captured by depth sensors [11] and object shapes are incomplete due to partial visibility or occlusions in the scene when using a single camera [3]. Therefore, depth completion methods are devised to fill in the missing depth values [12], but the pointcloud is only filled in on the visible side of the object and a depth camera is still required. As RGB-only cameras are more commonly available than RGB-D cameras, and less expensive, some works use wide-baseline stereo RGB cameras to perform human-to-robot handovers of container-like objects and address partial visibility and occlusions of the object [4, 6]. However, these works are limited to only container-like objects (rotationally symmetric shape), also assuming the object is held upright during the handover due to the presence of a content.
@@
Figure 1: We reconstruct the hand-object pointcloud from stereo RGB input for human-to-robot handover. (1) A safe grasp is selected for the handover and the robot moves in to grasp the object. (2) The object is delivered to a target location on the table. (3) The robot returns to its starting position.
@@
We propose StereoHO, a hand-object reconstruction method for wide-baseline stereo inputs. Our method quantifies the uncertainty of single-view predictions, e.g. caused by self-occlusions, by predicting low-dimensional embeddings of hand and object shapes as a probability distribution over learned codebooks of discrete shape embeddings. Individual predictions from each camera views are aggregated and decoded jointly to form a coherent stereo reconstruction. We train our model on synthetic 3D hand-object data [13] with a large amount of objects and without specific assumptions about the object’s geometry to ensure its generalisation ability. We integrate StereoHO in the first human-to-robot handover pipeline that jointly reconstructs hand and object shapes from stereo RGB inputs while tracking the hand over time. The pipeline also ensures multi-view consistency and reconstruction quality for accurate robot grasp predictions. Experiments [^1] Code and videos of the experiments available at https://qm-ipalab.github.io/StereoHO/ show that StereoHO outperforms previous methods in single-view and stereo settings for object reconstruction, while remaining competitive for hand reconstruction in terms of Chamfer distance. Furthermore, our pipeline can successfully handover objects of diverse shapes and appearances, including filled/empty transparent containers, boxes and thin objects.
@@ II Related work
Recent works in human-to-robot handover  [3, 8, 9, 10, 4, 6, 7] focused on markerless settings that are more applicable to daily human-to-robot interactions. Human grasps were classified into different types to select the corresponding predefined safe robot grasp direction and enable the robot to receive the object [7]. However, this approach does not consider diverse object shapes and limits the grasp types that the robot can handle. Later works improved the perception by capturing the shape of the hand and the object [3, 8, 9, 10, 4, 6], and used existing grasp estimation methods to predict possible grasps on the object shape. Grasps that would cause a collision between the robot and the human are removed for safety. A two-stage system was proposed to perform human-to-robot handovers of everyday objects [3]. The first stage segments the person and the object from the input RGB image. The depth image is lifted into a pointcloud and the background is removed using the segmentation masks. The second stage adds an imaginary plane in the grasping direction to approximate the task as a top-down grasp estimation problem using GG-CNN [14]. Using 6-DoF grasps allows more freedom in the robot’s movement, enabling more natural handovers [8].
@@
Existing methods for human-to-robot handovers [3, 8, 9, 10] and grasp estimation [14, 15], rely on pointcloud as a representation for hand and object shape. However, pointclouds of objects are view-dependent and are incomplete in the presence of occlusions [3] or when the objects are transparent and reflective [11]. Incomplete object shapes or missing pointclouds for transparent objects (Fig. 2) can results in inaccurate grasp estimations. To overcome this issue, depth restoration is a method to complete the pointcloud of transparent objects [11]. However, a depth camera is required which is not as commonly available as RGB cameras. Moreover, the completed pointcloud often only consists of the visible side of the object [12]. An alternative representation for object shapes is Signed Distance Dields (SDFs) that encode into a vector the signed distance between a set of points in 3D and their closest point on the 3D object [16]. As an implicit representation of smooth and continuous surfaces, SDFs distinguish the inside and outside parts of an arbitrary object and are easier to learn compared to pointclouds [16]. None of the existing works has used SDFs for reconstructing object-only or hand-object shapes for more accurate human-to-robot handovers.
@@
Other methods use stereo RGB images as input to reconstruct the shape of the object [4, 6]. The shape of containers, such as drinking glasses and cups, was reconstructed from stereo RGB images by assuming a rotationally symmetric geometry (Fig. 2) and exploiting the predicted segmentation masks of the object [17]. This enables human-to-robot handover by grasping at the centroid of the object [4]. Human safety is considered by estimating the 3D hand keypoints [6]. Robot grasps are sampled along the vertical axis of the object and grasp points that are too close to the hand keypoints are removed. However, these methods do not estimate 6-DoF grasps and thus limit the naturalness of the robot movement as the object must maintain an upright orientation during the handover. The assumption on the object shape is easily invalidated by non-container objects, causing failure in the object shape estimation.
@@
Figure 2: Limitations of existing human-to-robot handover approaches. Depth-based sensing fails on transparent objects (left). Relying on shape priors does not generalise (right).
@@
Joint hand-object reconstruction from RGB images in an interaction scenario is challenging due to severe mutual occlusions between the hand and the object [13]. Large hand-object interaction datasets, both synthetic and collected in real environments [13, 18, 19], have helped improve the reconstruction quality. Existing methods can be grouped into single-view and multi-view, and multi-view methods can be further split into dense and sparse. Single-view reconstruction methods [13, 20, 21, 22] learn 3D shape priors from hand-object datasets, but completing the shape of the unobserved parts of an object remains challenging. Multi-view methods [23, 24, 25, 26, 27] overcome the problems of occlusion and unobserved object parts by increasing the number of input views. Dense multi-view methods [23, 24, 25, 26] use a large number of input views to produce high quality reconstructions but are computationally expensive, limiting the applicability in reactive scenarios. Sparse multi-view methods [27] balances the reconstruction quality of dense methods with the speed of single-view methods. However, training only on synthetic data can affect the model’s ability to generalise to real world scenarios (sim-to-real gap) [27].
@@ III Proposed method
In this section, we introduce StereoHO for stereo hand-object reconstruction and our human-to-robot handover pipeline based on StereoHO.
@@ III-A Stereo-based RGB hand-object reconstruction
We design StereoHO to predict the joint surface of the hand ($H$) and object ($O$), represented as pointcloud $\mathcal{P}=[\mathcal{P}_{H},\mathcal{P}_{O}]$, from the RGB images of a stereo-based camera. Images are cropped and centred on the person’s hand holding an object. We formulate stereo-based hand-object shape reconstruction as a view-dependant image-to-shape classification task (encoding) followed by an aggregation step [20] and a transformation of the aggregated shape codes into the final joint hand-object pointcloud (decoding). Fig. 3 illustrates the inference pipeline of StereoHO. To avoid erroneously including background clutter in the reconstructed shape, we include the segmentation masks of the object, $M_{O}$, and the hand, $M_{H}$, as input to the model. We also predict the wrist pose $T_{H}$ for each view to relate the 3D shapes with the cameras views, assuming the camera poses are not available.
For the encoding part, we define a 3D grid centred on the hand wrist $T_{Hv}$, where $v\in\{L,R\}$ is the camera view. For each view $v$, we project the 3D position of each voxel into the image space using the wrist pose and associate each voxel with an image embedding, extracted with a convolutional neural network (e.g., ResNet-18 [28]), at the corresponding pixel location [20]. We use 3D convolutional layers to transform the resulting 4D tensor into a probability distribution $P_{v}\in[0,1]^{D\times D\times D\times C}$ with respect to a codebook of $C$ shape embeddings, $\mathcal{C}=\{e_{c}\in\mathbb{R}^{S}\}_{c=1}^{C}$. We combine the predicted probability distributions from each view into a single coherent prediction via element-wise multiplication: $P=P_{L}P_{R}$. We thus define another 4D tensor of shape embeddings that are selected from $\mathcal{C}$ based on the index of the highest probability for each element of the grid, $E\in\mathbb{R}^{D\times D\times D\times S}$. The same procedure is performed independently for the hand and the object, resulting in the two shape-based tensors, $E_{O}$ and $E_{H}$, respectively.
@@ III-B Learning discrete 3D shape embeddings
where $sg$ is the stop gradient operator. The hyper-parameter $\beta$ controls the scale of the commitment loss (last term), that helps stabilize the training [29]. We initialise the codebook with a uniform distribution. After training the models for object and hand, we only keep the trained SDF decoders and codebooks and link them to the image-to-shape encoder.
@@ III-C Training of the image-to-shape encoder
where $p_{gt}$ and $w_{c}$ are the ground-truth probability and weight for index $c$. We set a lower weight for the index $c$ corresponding to empty space as the majority of the space surrounding the hand and object in the 3D grid is empty.
@@ III-D Handover with stereo hand-object reconstruction
We integrate StereoHO into a pipeline for human-to-robot handovers (Fig. 4) and assume that i) the handover task starts with the person already holding the object and ii) two cameras are placed on either side of the robot, maximising the visible area of the hand and object.
For each frame of the approaching phase, StereoHO reconstructs the pointclouds of the hand and the object using image crops, pose of the wrist, and segmentation masks from each view. Therefore, for each camera view and for each frame, a hand-object detector [32] estimates the bounding boxes of both the hand and the object, $b_{H},b_{O}\in\mathbb{R}^{4}$. We use Fast Segment Anything [33] to segment the object ($M_{O}$) within the object bounding box [33] and we crop the image around the hand using $b_{H}$. From the image crop, we use FrankMocap [34] to predict the pose of the wrist, $T_{H}$ and the segmentation mask of the hand, $M_{H}$, after extracting the silhouette from the hand shape. We use both segmentation masks to remove the background from the image crop. We triangulate the centroid of the object mask and verify the reprojection error in each view is less than 5 pixels. Note that the pipeline does not perform the handover (failure case) if, within 10 s, no hand-object pair is detected in either of the views or the triangulation is not validated. Since the reconstruction quality can vary across frames, significantly affecting the estimated robot grasp, we compute the Intersection over Union (IoU) of the convex hull of the pointclouds with the segmentation masks, respectively. We update the pointclouds if there is an improvement in reconstruction quality (i.e., IoU $>$ IoU).
To enable the robot to receive the object, we use 6-DoF GraspNet [15] to estimate a set of $N$ candidate grasps, $\mathcal{G}=\{g_{n}\in SE(3)\}_{n=1}^{N}$, on the object pointcloud $\mathcal{P}^{\prime}_{O}$. To ensure human safety, we construct a 3D bounding box of the size of the gripper around each grasp, and we filter out any grasps with points from the hand point cloud $\mathcal{P}^{\prime}_{H}$ in its bounding box. Using the predicted wrist pose in the robot coordinate frame, $T_{HB}=T_{WB}T_{HW}$ where $T_{WB}$ is the robot base transform, we transform the predicted grasps from the hand-wrist to the robot coordinate frame. For each timestep, the robot control selects the grasp $g$ closest to the current gripper 6-DoF pose. We define a standoff position 15 cm away from the grasp point in the direction of the grasp. Once the robot reaches the standoff position, the robot enters a close-loop grasping stage and closes its fingers when the gripper reaches the selected grasp. After grasping, the robot retracts to the standoff position, delivers the object at a predefined location on the table, and returns to its initial position.

### v3

@@ Abstract
Jointly estimating hand and object shape facilitates the grasping task in human-to-robot handovers. Relying on hand-crafted prior knowledge about the geometric structure of the object fails when generalising to unseen objects, and depth sensors fail to detect transparent objects such as drinking glasses. In this work, we propose a method for hand-object reconstruction that combines single-view reconstructions probabilistically to form a coherent stereo reconstruction. We learn 3D shape priors from a large synthetic hand-object dataset, and use RGB inputs to better capture transparent objects. We show that our method reduces the object Chamfer distance compared to existing RGB based hand-object reconstruction methods on single view and stereo settings. We process the reconstructed hand-object shape with a projection-based outlier removal step and use the output to guide a human-to-robot handover pipeline with wide-baseline stereo RGB cameras. Our hand-object reconstruction enables a robot to successfully receive a diverse range of household objects from the human.
Index Terms:
Human-robot collaboration, physical human-robot interaction, deep learning for visual perception, perception for grasping and manipulation
@@ I Introduction
As humans and robots share the same workspace [1], exchanging objects is an essential skill to enable smooth collaboration. We consider the scenario where a human hands various everyday objects over to a robot, which aims to grasp the object while avoiding the human (human-to-robot handover [2, 3, 4]) and delivers the object at a predefined location (see Fig. 1). Objects can vary in their physical properties, such as appearance, transparency, shape, size, and mass [5]. Container-like objects, such as cups, drinking glasses, or boxes, can be empty or filled with content, affecting the perception of the appearance or the weight of the object [4, 6]. Reconstructing the hand and object shapes is essential in this highly interactive scenario. The object shape informs the robot on how to grasp the object and the hand shape allows the robot to avoid dangerous contact with the human.
@@
Visual modalities (RGB and depth images) are used in hand-object shape reconstruction for human-to-robot handover. Previous works [7, 3, 8, 9, 10] used depth sensors to capture the pointcloud of the scene and applied segmentation models to extract the pointcloud of the hand and object. However, transparent objects are difficult to detect with depth sensors [11] and object shapes are incomplete due to partial visibility or occlusions when using a single camera [3]. Therefore, depth completion methods are devised to fill in the missing depth values [12], but the pointcloud is only filled in on the visible side of the object and a depth camera is still required. As RGB-only cameras are more commonly available and less expensive than RGB-D cameras, some works [4, 6] used wide-baseline stereo RGB cameras to address partial visibility and occlusions of the object. However, these methods are limited to only container-like objects (rotationally symmetric shape) and assume that the object is held upright during the handover due to the presence of a content.
@@
Figure 1: We reconstruct the hand-object pointcloud from stereo RGB input for human-to-robot handover. (1) A safe grasp is selected for the handover and the robot moves in to grasp the object. (2) The object is delivered to a target location. (3) The robot returns to its home position.
To address the above challenges, in this paper we propose StereoHO [^1] Code and videos of the experiments are available at
https://qm-ipalab.github.io/StereoHO/
@@
, a hand-object reconstruction method for wide-baseline stereo inputs. Our method quantifies the uncertainty of single-view predictions, e.g. caused by self-occlusions, by predicting low-dimensional embeddings of hand and object shapes as a probability distribution over learned codebooks of discrete shape embeddings. Individual predictions from each camera views are aggregated and decoded jointly to form a coherent stereo reconstruction. We train our model on synthetic 3D hand-object data [13] with a large amount of objects and without specific assumptions about the object’s geometry to ensure its generalisation ability. We integrate StereoHO in the first human-to-robot handover pipeline that jointly reconstructs hand and object shapes from stereo RGB inputs while tracking the hand over time. The pipeline also ensures multi-view consistency and reconstruction quality for accurate robot grasp predictions. Experiments show that StereoHO outperforms previous methods in single-view and stereo settings for object reconstruction, while remaining competitive for hand reconstruction in terms of Chamfer distance. Furthermore, our pipeline can successfully account for objects of diverse shapes and appearances, including filled/empty transparent containers, boxes and thin objects.
@@ II Related work
Previous works [3, 8, 9, 10, 4, 6, 7] in human-to-robot handover focused on marker-less settings that are more applicable to daily human-to-robot interactions. Human grasps were classified into different types to select the corresponding predefined safe robot grasp direction and enable the robot to receive the object [7]. However, this approach does not consider diverse object shapes and limits the grasp types that the robot can handle. Capturing the shape of the hand and the object improves the perception, and using available grasp estimation methods allows to predict possible grasps on the object shape [3, 8, 9, 10, 4, 6]. Grasps that would cause a collision between the robot and the human are removed for safety. A two-stage system was proposed to perform human-to-robot handovers of everyday objects [3]. The first stage segments the person and the object from the input RGB image. The depth image is lifted into a pointcloud and the background is removed using the segmentation masks. The second stage adds an imaginary plane in the grasping direction to approximate the task as a top-down grasp estimation problem using GG-CNN [14]. Using 6-DoF grasps allows more freedom in the robot’s movement, enabling more natural handovers [8].
@@
Methods for human-to-robot handovers [3, 8, 9, 10] and grasp estimation [14, 15], rely on pointcloud as a representation for hand and object shape. However, pointclouds of objects are view-dependent and are incomplete in the presence of occlusions [3] or when the objects are transparent and reflective [11]. Incomplete object shapes or missing pointclouds for transparent objects (see Fig. 2) can results in inaccurate grasp estimations. To overcome this issue, depth restoration is a method to complete the pointcloud of transparent objects [11]. However, a depth camera is required which is not as commonly available as RGB cameras. Moreover, the completed pointcloud often only consists of the visible side of the object [12]. An alternative representation for object shapes is Signed Distance Dields (SDFs) that encode into a vector the signed distance between a set of points in 3D and their closest point on the 3D object [16]. As an implicit representation of smooth and continuous surfaces, SDFs distinguish the inside and outside parts of an arbitrary object and are easier to learn compared to pointclouds [16]. None of the existing works has used SDFs for reconstructing object-only or hand-object shapes for more accurate human-to-robot handovers.
@@
Other methods use stereo RGB images as input to reconstruct the shape of the object [4, 6]. The shape of containers, such as drinking glasses and cups, was reconstructed from stereo RGB images by assuming a rotationally symmetric geometry (see Fig. 2) and exploiting the predicted segmentation masks of the object [17]. This enables human-to-robot handover by grasping at the centroid of the object [4]. Human safety is considered by estimating the 3D hand keypoints [6]. Robot grasps are sampled along the vertical axis of the object and grasp points that are too close to the hand keypoints are removed. However, these methods do not estimate 6-DoF grasps and thus limit the naturalness of the robot movement as the object must maintain an upright orientation during the handover. The assumption on the object shape is easily invalidated by non-container objects, causing failure in the object shape estimation.
@@
Figure 2: Limitations of existing human-to-robot handover approaches. (a) Depth-based sensing fails on transparent objects. (b) Relying on shape priors does not generalise.
@@
Joint hand-object reconstruction from RGB images in an interaction scenario is challenging due to severe mutual occlusions between the hand and the object [13]. Large hand-object interaction datasets, both synthetic and collected in real environments [13, 18, 19], have helped improve the reconstruction quality. Methods for hand-object reconstruction can be grouped into single-view and multi-view, and multi-view methods can be further split into dense and sparse. Single-view reconstruction methods [13, 20, 21, 22] learn 3D shape priors from hand-object datasets. However, completing the shape of the unobserved parts of an object remains challenging. Multi-view methods [23, 24, 25, 26, 27] overcome the problems of occlusion and unobserved object parts by increasing the number of input views. Dense multi-view methods [23, 24, 25, 26] use a large number of input views to produce high quality reconstructions but are computationally expensive, limiting the applicability in reactive scenarios. Sparse multi-view methods [27] balances the reconstruction quality of dense methods with the speed of single-view methods. However, training only on synthetic data can affect the model’s ability to generalise to real world scenarios (sim-to-real gap) [27].
@@ III Proposed method
In this section, we introduce StereoHO for stereo hand-object reconstruction and present our human-to-robot handover pipeline based on StereoHO.
@@ III-A Stereo-based RGB hand-object reconstruction
We design StereoHO to predict the joint surface of the hand ($H$) and object ($O$), represented as pointcloud ($\mathcal{P}=[\mathcal{P}_{H},\mathcal{P}_{O}]$), from the RGB images of a stereo-based camera. Images are cropped and centred on the person’s hand holding an object. We formulate stereo-based hand-object shape reconstruction as a view-dependant image-to-shape classification task (encoding) followed by an aggregation step [20] and a transformation of the aggregated shape codes into the final joint hand-object pointcloud (decoding). Fig. 3 illustrates the inference pipeline of StereoHO. To avoid erroneously including background clutter in the reconstructed shape, we include the segmentation masks of the object ($M_{O}$) and the hand ($M_{H}$) as input to the model. We also predict the wrist pose ($T_{H}$) for each view to relate the 3D shapes with the cameras views, assuming the camera poses are not available.
For the encoding part, we define a 3D grid centred on the hand wrist $T_{Hv}$, where $v\in\{L,R\}$ is the camera view. For each view $v$, we project the 3D position of each voxel into the image space using the wrist pose and associate each voxel with an image embedding, extracted with a convolutional neural network (e.g., ResNet-18 [28]), at the corresponding pixel location [20]. We use 3D convolutional layers to transform the resulting 4D tensor into a probability distribution $P_{v}\in[0,1]^{D\times D\times D\times C}$ with respect to a codebook of $C$ shape embeddings, $\mathcal{C}=\{e_{c}\in\mathbb{R}^{S}\}_{c=1}^{C}$.
We combine the predicted probability distributions from each view into a single coherent prediction via element-wise multiplication: $P=P_{L}P_{R}$. We thus define another 4D tensor of shape embeddings that are selected from $\mathcal{C}$ based on the index of the highest probability for each element of the grid, $E\in\mathbb{R}^{D\times D\times D\times S}$. The same procedure is performed independently for the hand and the object, resulting in the two shape-based tensors, $E_{O}$ and $E_{H}$, respectively.
@@ III-B Learning discrete 3D shape embeddings
where $sg[\cdot]$ is the stop gradient operator. The hyper-parameter $\beta$ controls the scale of the commitment loss (last term), that helps stabilize the training [29]. We initialise the codebook with a uniform distribution. After training the models for object and hand, we only keep the trained SDF decoders and codebooks and link them to the image-to-shape encoder.
@@ III-C Training of the image-to-shape encoder
where $p_{gt}$ and $w_{c}$ are the ground-truth probability and weight for index $c$. We set a lower weight for the index $c$ corresponding to empty space, as the majority of the surroundings of the hand and object in the 3D grid is empty.
@@ III-D Handover with stereo hand-object reconstruction
We integrate StereoHO into a pipeline for human-to-robot handovers (see Fig. 4) and assume that i) the handover task starts with the person already holding the object and ii) two cameras are placed on each side of the robot, maximising the visible area of the hand and object.
For each frame of the approaching phase, StereoHO reconstructs the pointclouds of the hand and the object using image crops, pose of the wrist, and segmentation masks from each view. Therefore, for each camera view and for each frame, a hand-object detector [32] estimates the bounding boxes of both the hand and the object, $b_{H},b_{O}\in\mathbb{R}^{4}$. We use Fast Segment Anything [33] to segment the object ($M_{O}$) within the object bounding box [33] and we crop the image around the hand using $b_{H}$. From the image crop, we use FrankMocap [34] to predict the pose of the wrist ($T_{H}$) and the segmentation mask of the hand ($M_{H}$) after extracting the silhouette from the hand shape. We use both segmentation masks to remove the background from the image crop. We triangulate the centroid of the object mask and verify that the reprojection error in each view is less than 5 pixels. Note that the pipeline does not perform the handover (failure case) if, within 10 s, no hand-object pair is detected in either of the views or the triangulation is not validated. Since the reconstruction quality can vary across frames, significantly affecting the estimated robot grasp, we compute the Intersection over Union (IoU) of the convex hull of the pointclouds with the segmentation masks, respectively. We update the pointclouds if there is an improvement in reconstruction quality (i.e., IoU $>$ IoU).
To enable the robot to receive the object, we use 6-DoF GraspNet [15] to estimate a set of $N$ candidate grasps, $\mathcal{G}=\{g_{n}\in SE(3)\}_{n=1}^{N}$, on the object pointcloud $\mathcal{P}^{\prime}_{O}$. To ensure human safety, we construct a 3D bounding box of the size of the gripper around each grasp, and we filter out any grasps with points from the hand point cloud $\mathcal{P}^{\prime}_{H}$ in its bounding box. Using the predicted wrist pose in the robot coordinate frame, $T_{HB}=T_{WB}T_{HW}$, where $T_{WB}$ is the robot base transform, we transform the predicted grasps from the hand-wrist to the robot coordinate frame. For each timestep, the robot control selects the grasp $g$ closest to the current gripper 6-DoF pose. We define a standoff position 15 cm away from the grasp point in the direction of the grasp. Once the robot reaches the standoff position, the robot enters a close-loop grasping stage and closes its fingers when the gripper reaches the selected grasp. After grasping, the robot retracts to the standoff position, delivers the object at a predefined location on the table, and returns to its initial position.